Week of 160627

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (Chair and minutes), Cedric Serfon (ATLAS), Maarten Litmaath (ALICE), Ken Bloom (CMS), Vladimir (LHCb), Julia Andreeva (WLCG), Jerome Belleman (Computing), Sebastien Masson (DB), Maria Dimou (GGUS), Andrea Manzi (MW Officer), Vincent Brillault (Security), Marian Babik (Monitoring and Networking)
  • remote: Francesco Noferini (CNAF), Rolf Rumler (IN2P3), Sang-Un Ahn (KISTI), Dimitri (KIT), Christian (NDGF), Andrew Pickford (NL-T1), John Kelly (RAL), Di Qing (TRIUMF).

Experiments round table:

  • ATLAS reports (raw view) -
    • LHC and data processing running full steam. Record lumi reached !
    • (Re)Testing T0 spillover (i.e. processing of T0 data on the Grid)
    • Problems:
      • Alarm ticket on Sunday due to 100% failure rate writing to Castor (GGUS:122325). Moreover GGUS:122208 (Alarm ticket from last week) still open : "timeout in an internal communication between the CASTOR gridftp server and the diskmanager daemon"
      • RRC-KI-T1 file loss follow-up : 37k unique files lost out of 131k lost files. Will be regenerated.
      • EOS namespace crashed and restarted (twice last week)
      • Slow transfers from AGLT2 to CERN : GGUS:122293. Will work on automatic identification of these slow links.

Maria will follow up with the CERN Storage team the open tickets. Regarding the slow transfers, Cedric adds that it would be good to get automatic information from perfSONAR. Marian agrees this is interesting to know but detecting slow links doesn't prevent from doing more detailed investigation to understand why slow throughput is happening. Marian asks to be added to the ticket so that he can also follow this up.

  • CMS reports (raw view) -
    • In general things are fine on the grid, we're keeping sites busy.
    • Quickly filling up T0 EOS space because we are not transferring data fast enough out to T1's. The right solution is under study, but in my opinion it is fixable.
    • Interesting glitch with xrootd federation over the weekend in which all of the OSG region sites got unsubscribed from the federation. Fixed but trying to understand what happened.

  • ALICE -

  • LHCb reports (raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction and user jobs on the Grid
    • Site Issues
      • T0:
        • Bad handling of checksum error in FTS: GGUS:122245 (needs development)
        • Wrong configuration of worker nodes: GGUS:122187 (in progress)
        • Problem with EOS: GGUS:122100
      • GRIDKA:
        • Unscheduled Downtime

Maria will follow up the EOS ticket, now in open state for a while, after the meeting.

Sites / Services round table:

  • ASGC: NA
  • BNL: NA
  • CNAF: The intervention on the frontier router done on 21th (18-20) was completed accordingly to the schedule (NTR). Today we closed the CMS queue to proceed with the intervention on a DDN which will take place tomorrow (down on 28th from 8.00 to 20.00).
  • FNAL: NTR
  • GridPP: NA
  • IN2P3: NTR
  • JINR: Nothing notable to report.
  • KISTI: NTR
  • KIT: Due to network maintenance, network was unreachable for 20min in the morning. Tape downtime starting tomorrow and lasting two days. No read or write possible. All experiments affected.
  • NDGF: dCache update on headnodes Wednesday 29/6. Atlas and Alice data affected. Should only be a ~5 min interruption during reboot.
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: NA
  • PIC: NA
  • RAL: No news for the tape controller SW, although this has little impact on operations.
  • TRIUMF: NTR

  • CERN computing services: NTR
  • CERN storage services: NA
  • CERN databases: CMS Online database was touched by performance issues on Thursday due to unforeseen load added to the database on Wednesday evening. Load had to be rebalanced between the DB nodes (one node still down due to HW issues). REMINDER: Please also do not use the online db unless you need to access them for online production operations. For all other uses case you should use ADG. Please contact the DB experts if you need to transfer production data to development.
  • GGUS: NTR
  • MW Officer:
    • No news on the Top Priority ticket GGUS:120586 reported by LHCb and assigned to CESNet. We reported about it during the EGI URT meeting, but no one from CESNet was connected to update about its status.
  • Security: NTR
  • Monitoring:
    • Creation of the monthly availability reports for May blocked by two recomputation requests ( OU_OCHEP_SWT2 PT-LIP-LCG-Tier2. ). Waiting for ATLAS to apply the correction
  • Network:
    • CERN: Upgrade of the LCG distribution router (Mon Jul 04 10:00:00 CEST 2016) followed by reboot

AOB:

Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2016-06-28 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback